# 1]Centre Étude et simulation du climat à l’échelle régionale (ESCER), Département des Sciences de la Terre et de l’Atmosphère, Université du Québec à Montréal, Montréal, Québec, Canada

### v2

@@ 1.4 Initial Condition Perturbations
Table 2 presents the maximum perturbations applied to the initial conditions, alongside the range of values sampled in the ensemble simulation. The range is computed by finding the max perturbation of all the ensemble members with respect to the mean. Overall, the applied perturbations during optimization remain within or below the variability represented in the stochastic ensemble.
Quantity Max Perturbation (N = 50) Max Perturbation (N = 75) Range of Perturbations Ensemble Run Surface Pressure (Pa) 0.69 0.47 0.0 Specific Humidity $2.34\times 10^{-3}$ $1.11\times 10^{-3}$ $3.20\times 10^{-3}$ Specific Cloud Ice Water Content (kg kg-1) $7.61\times 10^{-6}$ $5.19\times 10^{-6}$ $4.35\times 10^{-5}$ Specific Cloud Liquid Water Content (kg kg-1) $1.50\times 10^{-5}$ $2.42\times 10^{-5}$ $7.61\times 10^{-5}$ Temperature (K) 4.83 4.99 7.60 U-component of windspeed (m s-1) 8.37 5.59 12.70 V-component of windspeed (m s-1) 4.77 4.12 7.94
Table 2: Maximum perturbations for different run sizes compared to the range of perturbations applied on the ensemble run by the stochastic model.
@@ 2.1 NeuralGCM temperature evaluation
We first evaluate the ability of the NeuralGCM model to simulate a temperature climate to ERA5. Figure 1a) presents the 6-hourly temperature distribution for a NeuralGCM simulation and the ERA5 data over the 1981-2020 40-year period, averaged within the domain of interest (highlighted in the blue box in Fig. 2). The NeuralGCM simulation closely approximates the ERA5 distribution, demonstrating its ability to reproduce key statistical characteristics of the temperature distribution in this region. We highlight the 95th and 99th percentile values for both the model and ERA5. We note that the model has slightly colder hot extremes than ERA5.
Figure 1: (a) Histograms of 6-hourly temperature values from a 40-year NeuralGCM simulation (green) and from ERA5 (orange) reanalysis over the study domain outlined in Figure 4. Dashed line indicates the 95th and 99th percentiles. (b) Time series of temperature forecasts at 1000hPa from NeuralGCM with 10-, 6-, 4-, and 2-day lead times (green colored lines) compared with ERA5 reanalysis data (orange line) for the PN2021 heatwave. Grey area highlights the targeted time range for the optimization process.
Next the ability of the NeuralGCM model to forecast the PN2021 heatwave against the ERA5 reanalysis data is evaluated. Figure 1b) shows NeuralGCM forecasted surface temperatures during the PN2021 heatwave for five lead times ranging from 10 to 2 days. At a 10-day lead time, the NeuralGCM predictions follow closely the ERA5 data until about day 8, where they deviate leading to a lack of heatwave and extreme temperatures. At an 8-day lead time, the simulation substantially enhances temperature during the heatwave period but still shows large underestimations of peaks intensities, with differences of about $6^{\circ}$. At 2-, 4-, and 6-day lead times, the NeuralGCM model captures well the general pattern of temperature variations shown by ERA5, including the occurrence of very high temperatures during the heatwave event. However, most forecasts underestimate the peak intensity compared to ERA5 by a few degrees Celsius, particularly during the days after the peak intensity.
@@ 2.2 Optimizing extreme temperatures
Figure 2: Top row: Evolution of the difference in 500-hPa geopotential height ($\Delta Z$, in km) between the optimized simulation and the control run. Black contours (optimized run) outline the amplified Rossby wave pattern, with deeper troughs and higher ridges compared to the control. Middle row: The difference in 1000-hPa temperature (C) between the optimized and control simulations. Bottom row: Th difference between the 500-hPa geopotential height and 1000-hPa temperature averaged over the target domain.
@@
Figure 3: Top row: cross section of 500 hPa geopotential along a fixed latitude. Bottom row: the amplitude difference of the Fourier spectrum, including wavenumbers 1–5 highlighted in grey. Red dots highlight wavenumbers 2-5.
The optimization process relies on gradient descent (see Sect. 1 for details), which requires choosing the number of gradient descent steps. Figure 4 shows hourly time series of $T_{1000}$ (Figure 4a) and $\phi_{500}$ (Figure 4c) for two optimized trajectories with $N=50$ and $N=75$ steps, alongside a 75-member ensemble and its mean, all initialized 8 days before the peak of the event. Notably, the $T_{1000}$ time series (Fig. 4 a) reveals that both optimized trajectories attain values beyond the range exhibited by any individual ensemble member. In other words, the proposed method allows us to find extreme temperature values that are more extreme than those found using a 75-member ensemble using only 50 iterations (a 33% reduction in computational cost). Specifically, the trajectory from 50 steps reaches a peak temperature of $37.0^{\circ}$, while the trajectory after 75 steps attains $38.9^{\circ}$. Compared to the mean of the ensemble, we reach anomalies of $14.0^{\circ}$. For the 500-hPa geopotential height (Fig. 4c), both optimized trajectories show similarly elevated values, once again exceeding the range spanned by the 75-member ensemble. Importantly, the trajectory from the 75-step optimized run maintains a more sustained increase of the 500-hPa geopotential height compared to that from the 50-step run. The measured intensity (Fig. 4e) and length of the event (Fig. 4f) are increased in both optimized runs compared to the ensemble mean. Both optimized runs produce a 6-day-long heatwave event, differing only in intensity, with the 75-step run having a $1.0^{\circ}$ higher intensity than the 50-step run. Notably, both optimized solutions exceed the intensities spanned within the 75-members ensemble. Figure 4b,d present the temperature and geopotential fields from the 75-iteration run. The temperature pattern features a maximum over the targeted region, albeit slightly to the south, while the 500 hPa geopotential height field exhibits an anticyclone directly overhead.
@@ 2.3 Sensitivity of other variables
Figure 5 shows optimized trajectories for near-surface wind (zonal U and meridional V components), specific humidity, and sea-level pressure. In the optimized run, the U wind lies consistently at the lower end of the ensemble spread during the event, and specific humidity likewise tracks near the lower end. Such concurrent reductions in near-surface wind speeds and humidity are consistent with the physical mechanisms that underlie heatwave intensification. In contrast, the V wind and sea-level pressure exhibit only slight positive anomaly above the ensemble mean. All optimized trajectories remain entirely within the bounds defined by the non-optimized ensemble members.
Figure 5: Time series of (a) surface pressure (b) near-surface specific humidity (c) U-component of wind and (d) V-component of wind at 1000 hPa. Data from the optimized trajectory are shown alongside the individual ensemble members (in grey) and the ensemble mean (thick black line).
@@ 2.4 Sensitivity to NeuralGCM resolution
To test how resolution affects our optimization, we reran the optimization problem on NeuralGCM at a finer $1.4^{\circ}$ resolution (Figure 6). In the unperturbed control, the high-resolution ensemble mean reduces the warm bias against ERA5 and more accurately captures the peak and decay of the PN2021 heatwave. Once optimized, the $1.4^{\circ}$ run again produces peak surface-temperature anomalies that exceed the 75-member ensemble maximum, and 500 hPa geopotential-height anomalies that surpass the control even more than the $2.8^{\circ}$ case. These enhanced temperature and geopotential height anomalies persist through the extended target period, demonstrating that the optimization delivers sustained extremes even at higher resolution.
@@ 3 Discussion
Our findings demonstrate that differentiable climate models, exemplified by NeuralGCM, offer a powerful tool for constructing extreme heatwave storylines through gradient-based optimization of initial conditions. By perturbing initial conditions, we identified alternative trajectories with slightly different synoptic-scale conditions that amplify the PN2021 heatwave intensity by $3.7^{\circ}$ according to NeuralGCM. These results align with prior studies linking extreme heat to persistent blocking patterns (Screen and Simmonds, 2014; Mass et al., 2024). Specifically, the optimized geopotential height anomalies and spectrum reflect enhanced blocking dynamics with an amplification of wavenumbers 1-5.
@@
While NeuralGCM resolves large-scale dynamics, its omission of land-atmosphere feedbacks (e.g., soil moisture (Duan et al., 2025)) likely results in a conservative estimate of heatwave amplification. For instance, soil moisture–temperature coupling is known to cause stronger heatwave persistence (e.g., (Suarez-Gutierrez et al., 2020; White et al., 2023)), implying that the model might underestimate extremes when such feedbacks are neglected. Additionally, the model’s coarse horizontal resolution ($2.8^{\circ}$) introduces biases in capturing localized extreme conditions associated with the PN2021 event. As show with the higher resolution simulation, the use of finer grids ($1.4^{\circ}$) allows for more accurate estimates of the extreme temperatures of the event although the main dynamical changes remain similar to the coarse resolution. In addition, while we have chosen to use NeuralGCM for this study, the method could be applied to any model which had automatic differentiation implemented. This includes all existing purely data-driven models, including GraphCast, Pangu-weather, FourCastNet, and FuXi. While data-driven models could provide faster predictions at higher resolution, their dual-initial-condition requirement introduces ambiguity about finding optimal initial conditions. For instance, Vonich and Hakim (2024) optimized both inputs for GraphCast to reconstruct the 2021 heatwave, but this approach demands simultaneous perturbation of two distinct states. Validation against models like Pangu demonstrated that results remained consistent despite this added complexity, suggesting robustness in the dual-input framework. However, NeuralGCM’s hybrid design simplifies the workflow by requiring only a single initial condition. In addition, Selz and Craig (2023) show that data-driven forecast models tend to poorly represent small perturbations, often filtering them out, which could impact the method. The extent to which this affects hybrid models, which has a dynamical core, such as NeuralGCM remains unclear.
@@
The optimization process involves making several decisions and setting specific parameters. While we do not present the results here, we have explored a limited subset of the broader hyperparameter space—specifically, the learning rate ($\alpha$) and the loss function parameters ($\beta$,$\vec{\lambda}$). We found that, for the loss function parameters defined in Sect. 1, a large learning rate induces instability, causing substantial perturbations without a corresponding increase in temperature and in some cases, triggering numerical instabilities that caused the simulation to fail. To ensure stability, $\alpha$ must be on the order of $10^{-9}$ or smaller. This stability condition varies in a nonlinear and nontrivial manner with changes in $\vec{\lambda}$, and a thorough exploration of the hyperparameter space lies beyond the scope of this study. Fine-tuning the parameters might allow for improved efficiency in computational cost, where a more extreme solution is found with a reduced number of steps.
@@
To evaluate the robustness and physical realism of the NeuralGCM-optimized initial conditions, these perturbations should be tested in a conventional numerical weather prediction model (e.g., Environment and Climate Change Canada’s model Buehner et al. (2015)). Such cross-model validation would reveal the universality of the results and help isolate NeuralGCM-specific biases. Additionally, running these scenarios in a fully physical model would explicitly account for land–atmosphere interactions and feedbacks, and assess whether the extreme trajectories persist under more detailed dynamics and physics.
@@
The computational efficiency of the ML and hybrid models coupled with their differentiable properties, opens avenues for exploring extreme events—from heatwaves to precipitation extremes and compound disasters. For example, a similar optimization problem could be formulated for StormCast (Pathak et al., 2024) to allow us to search for extreme precipitation events in an emulated regional climate model. One could also formulate the loss function such that the large deviation theory rate function is minimized, leading to “typical” trajectories of extremes (Grafke and Vanden-Eijnden, 2019; Zakine and Vanden-Eijnden, 2023). We could also envision loss functions with hard constraints on the perturbation which impose conservation laws as opposed to simply imposing small perturbations.
@@
We introduce a differentiable-storyline framework that leverages automatic differentiation in hybrid climate models to directly optimize initial conditions and generate physically coherent extreme-heatwave trajectories at a fraction of the computational cost of traditional ensemble methods. When applied to the PN2021 heatwave, our approach produces intensifications of nearly $3.7^{\circ}$ by isolating high-impact circulation patterns, specifically enhanced blocking and Rossby-wave, demonstrating both its dynamical fidelity and efficiency. While this proof-of-concept focuses on NeuralGCM and a single case study, the optimization paradigm is agnostic to model architecture and event type, offering a transformative tool for rapid, process-based risk assessment of diverse climate extremes in a warming world.
@@
Data availability: All data generated using the code and ERA5 data used in this study is openly available from the Copernicus Data Store (https://doi.org/10.24381/cds.143582cf) Hersbach et al. (2020).

### v3

@@ 1.3 Initial Condition Perturbations
Table 2 presents the maximum perturbations applied to the initial conditions, alongside the range of values sampled in the ensemble simulation. The range is computed by finding the max perturbation of all the ensemble members with respect to the mean. Overall, the applied perturbations during optimization remain within or below the variability represented in the stochastic ensemble. Visualizing the perturbations directly is challenging due to their high dimensionality, but their spatial spectrum provides some insight. We present the perturbation spectrum in App. B.
Quantity / Experiment EXP50 EXP75 75-member ensemble #\# of steps 50 75 - Surface Pressure (Pa) 0.69 0.47 0.0 Specific Humidity $2.34\times 10^{-3}$ $1.11\times 10^{-3}$ $3.20\times 10^{-3}$ Specific Cloud Ice Water Content (kg kg-1) $7.61\times 10^{-6}$ $5.19\times 10^{-6}$ $4.35\times 10^{-5}$ Specific Cloud Liquid Water Content (kg kg-1) $1.50\times 10^{-5}$ $2.42\times 10^{-5}$ $7.61\times 10^{-5}$ Temperature (K) 4.83 4.99 7.60 U-component of windspeed (m s-1) 8.37 5.59 12.70 V-component of windspeed (m s-1) 4.77 4.12 7.94
Table 2: Maximum perturbations over the full 3D fields for different run sizes compared to the range of perturbations applied on the ensemble run by the stochastic model.
@@ 2.1 NeuralGCM temperature evaluation
We first evaluate the ability of the NeuralGCM2.8 model to simulate a temperature summer climate (June, July, August) compared to ERA5. Figure 1a) presents the 6-hourly temperature distribution for a NeuralGCM2.8 simulation and the ERA5 data over the 1981-2020 40-year period, averaged within the domain of interest (highlighted in the blue box in Fig. 4). The NeuralGCM2.8 simulation closely approximates the ERA5 distribution, demonstrating its ability to reproduce key statistical characteristics of the temperature distribution in this region. We highlight the 95th and 99th percentile values for both the model and ERA5. We note that the model has slightly colder hot extremes than ERA5.
Figure 1: (a) Histograms of 6-hourly temperature values from a 40-year NeuralGCM simulation (green) and from ERA5 (orange) reanalysis over the study domain outlined in Figure 4. Dashed line indicates the 95th and 99th percentiles. (b) Time series of temperature forecasts at 1000-hPa from NeuralGCM with 10-, 6-, 4-, and 2-day lead times (green colored lines) compared with ERA5 reanalysis data (orange line) for the PN2021 heatwave. Grey area highlights the targeted time range for the optimization process.
Next the ability of the NeuralGCM2.8 model to forecast the PN2021 heatwave against the ERA5 reanalysis data is evaluated. Figure 1b) shows NeuralGCM2.8 forecasted surface temperatures during the PN2021 heatwave for five lead times ranging from 10 to 2 days. At a 10-day lead time, the NeuralGCM2.8 predictions follow closely the ERA5 data until about day 8, where they deviate leading to a lack of heatwave and extreme temperatures. At an 8-day lead time, the simulation substantially enhances temperature during the heatwave period but still shows large underestimations of peaks intensities, with differences of about $6^{\circ}$C. At 2-, 4-, and 6-day lead times, the NeuralGCM2.8 model captures well the general pattern of temperature variations shown by ERA5, including the occurrence of very high temperatures during the heatwave event. However, most forecasts underestimate the peak magnitude compared to ERA5 by a few degrees Celsius, particularly during the days after the peak of the event.
This underestimation of the extreme heat, to our knowledge, is due to two factors: 1) there seems to be a dependence on capturing the extreme with the coarseness of the model, when we increase the resolution to the $1.4^{\circ}$ model, the prediction quality improves (see Fig. 1C)). 2) other studies have evaluated the ability of simulating extreme heatwave storylines and found that the model lacking processes, such land-surface feedbacks led to under representation of extreme (duan2024aibasedclimatemodelevaluation).
@@ 2.2 Optimizing extreme temperatures
Figure 2: Top row: Evolution of the difference in 500-hPa geopotential height ($\Delta Z$, in km) between the optimized simulation and the control run for EXP75. Black contours (optimized run) outline the amplified Rossby wave pattern, with deeper troughs and higher ridges compared to the control. Middle row: The difference in 1000-hPa temperature (C) between the optimized and control simulations. Bottom row: Th difference between the 500-hPa geopotential height and 1000-hPa temperature averaged over the target domain.
@@
Figure 3: Top row: cross section of 500-hPa geopotential along a fixed latitude for EXP75. Bottom row: the amplitude difference of the Fourier spectrum, including wavenumbers 1–5 highlighted in grey. Red dots highlight wavenumbers 2-5.
The optimization process relies on gradient descent (see Sect. 1 for details), which requires choosing the number of gradient descent steps. Figure 4 shows hourly time series of $T_{1000}$ (Figure 4a) and $\phi_{500}$ (Figure 4c) for two optimized trajectories with $N=50$ (EXP50) and $N=75$ (EXP75) steps, alongside a 75-member ensemble and its mean, all initialized 8 days before the peak of the event. Notably, the $T_{1000}$ time series (Fig. 4 a) reveals that both optimized trajectories attain values beyond the range exhibited by any individual ensemble member. In other words, the proposed method allows us to find extreme temperature values that are more extreme than those found using a 75-member ensemble using only 50 iterations (a 33% reduction in computational cost relative to generating the 75-member ensemble, calculated as (7550)/75(75-50)/75). Notably, this more efficient 50-step optimization run produces a trajectory more extreme than any member of the 75-member ensemble. Specifically, the trajectory from 50 steps reaches a peak temperature of $37.0^{\circ}$C, while the trajectory after 75 steps attains $38.9^{\circ}$C. Compared to the mean of the ensemble, we reach anomalies of $14.0^{\circ}$C. For the 500-hPa geopotential height (Fig. 4c), both optimized trajectories show similarly elevated values, once again exceeding the range spanned by the 75-member ensemble. Importantly, the trajectory from the 75-step optimized run maintains a more sustained increase of the 500-hPa geopotential height compared to that from the 50-step run. The measured intensity (Fig. 4e) and length of the event (Fig. 4f) are increased in both optimized runs compared to the ensemble mean. Both optimized runs produce a 6-day-long heatwave event, differing only in intensity, with the 75-step run having a $1.0^{\circ}$C higher intensity than the 50-step run. Notably, both optimized solutions exceed the intensities spanned within the 75-members ensemble. Figure 4b,d present the temperature and geopotential fields from the 75-iteration run. The temperature pattern features a maximum over the targeted region, albeit slightly to the south, while the 500-hPa geopotential height field exhibits an anticyclone directly overhead. The resulting fields from the optimized solution are consistent with what is seen in ERA5 data for the PN2021 event as can be seen in App. A.
@@ 2.3 Sensitivity of other variables
Figure 5 shows optimized trajectories for near-surface wind (zonal U and meridional V components), specific humidity, surface pressure, and temperature advection. In the optimized run, the U wind lies consistently at the lower end of the ensemble spread during the event, and specific humidity likewise tracks near the lower end. Such concurrent reductions in near-surface wind speeds and humidity are consistent with the physical mechanisms that underlie heatwave intensification. In contrast, the V wind and sea-level pressure exhibit only slight positive anomaly above the ensemble mean. All optimized trajectories remain entirely within the bounds defined by the non-optimized ensemble members. While the variables are within range of the ensemble envelope (i.e., not extreme), there might be a confluence of factors that leads to the extreme.
Figure 5: Time series of (a) surface pressure, (b) near-surface specific humidity, (c) U-component of wind, (d) V-component of wind, (e) and (f) temperature advection for each component at 1000-hPa . Data from the optimized trajectory are shown alongside the individual ensemble members (in grey) and the ensemble mean (thick black line).
@@ 2.4 Sensitivity to NeuralGCM resolution
To test how resolution affects our optimization, we reran the optimization problem on NeuralGCM at a finer $1.4^{\circ}$ resolution (Figure 6). The same parameters as in Table. 1 is used for this set of experiments. In the unperturbed control, the high-resolution ensemble mean reduces the warm bias against ERA5 and more accurately captures the peak and decay of the PN2021 heatwave. Once optimized, the $1.4^{\circ}$ run again produces peak surface‐temperature anomalies that exceed the 75‐member ensemble maximum, and 500-hPa geopotential‐height anomalies that surpass the control even more than the $2.8^{\circ}$ case. These enhanced temperature and geopotential height anomalies persist through the extended target period, demonstrating that the optimization delivers sustained extremes even at higher resolution.
@@ 3 Discussion
Our findings demonstrate that differentiable climate models, exemplified by NeuralGCM, offer a powerful tool for constructing extreme heatwave storylines through gradient-based optimization of initial conditions. By perturbing initial conditions, we identified alternative trajectories with slightly different synoptic-scale conditions that amplify the PN2021 heatwave intensity by $3.7^{\circ}$C according to NeuralGCM. These results align with prior studies linking extreme heat to persistent blocking patterns (Screen2014; ThePacificNorthwestHeatWave). Specifically, the optimized geopotential height anomalies and spectrum reflect enhanced blocking dynamics with an amplification of wavenumbers 1-5. The resulting temperature and geopoentatial increase has realistic features when comparing the to ERA5 data as can be seen in App. A.
@@
While NeuralGCM resolves large-scale dynamics, its omission of land-atmosphere feedbacks (e.g., soil moisture (duan2024aibasedclimatemodelevaluation)) likely results in a conservative estimate of heatwave amplification. For instance, soil moisture–temperature coupling is known to cause stronger heatwave persistence (e.g., (Suarez-Gutierrez2020; White2023)), implying that the model might underestimate extremes when such feedbacks are neglected. Additionally, the model’s coarse horizontal resolution ($2.8^{\circ}$) introduces biases in capturing localized extreme conditions associated with the PN2021 event. As show with the higher resolution simulation, the use of finer grids ($1.4^{\circ}$) allows for more accurate estimates of the extreme temperatures of the event although the main dynamical changes remain similar to the coarse resolution. In addition, while we have chosen to use NeuralGCM for this study, the method could be applied to any model which had automatic differentiation implemented. This includes all existing purely data-driven models, including GraphCast, Pangu-weather, FourCastNet, and FuXi. While data-driven models could provide faster predictions at higher resolution, their dual-initial-condition requirement introduces ambiguity about finding optimal initial conditions. For instance, vonich2024predictabilitylimit2021pacific optimized both inputs for GraphCast to reconstruct the 2021 heatwave, but this approach demands simultaneous perturbation of two distinct states. Validation against models like Pangu demonstrated that results remained consistent despite this added complexity, suggesting robustness in the dual-input framework. However, NeuralGCM’s hybrid design simplifies the workflow by requiring only a single initial condition. In addition, https://doi.org/10.1029/2023GL105747 show that data-driven forecast models tend to poorly represent small perturbations, often filtering them out, which could impact the method. The extent to which this affects hybrid models, which has a dynamical core, such as NeuralGCM remains unclear.
@@
The optimization process involves making several decisions and setting specific parameters. While we do not present the results here, we have explored a limited subset of the broader hyperparameter space—specifically, the learning rate ($\alpha$) and the loss function parameters ($\beta$$\vec{\lambda}$). We found that, for the loss function parameters defined in Sect. 1, a large learning rate induces instability, causing substantial perturbations without a corresponding increase in temperature and in some cases, triggering numerical instabilities that caused the simulation to fail. To ensure stability, $\alpha$ must be on the order of $10^{-9}$ or smaller. This stability condition varies in a nonlinear and nontrivial manner with changes in $\vec{\lambda}$. Furthermore, the consistency of the results across the EXP50 and EXP75 experiments and the simulations at two different resolutions—all of which yield trajectories more extreme than the 75-member stochastic ensemble—suggests that the optimized perturbations are not simply initialization artifacts. However, a systematic quantification of the sensitivity to the initial state and a thorough exploration of the hyperparameter space lie beyond the scope of this study. Fine-tuning the parameters might allow for improved efficiency in computational cost, where a more extreme solution is found with a reduced number of steps. In addition we note that we have chosen to optimize the 1000-hPa level but, NeuralGCM utilizes $\sigma$-coordinates. Over regions with significant elevation, such as the Canadian Rockies, the 1000-hPa geopotential surface is often below ground level. Using the 1000-hPa temperature ($T_{1000}$) can therefore yield physically inconsistent values when optimizing for near-surface extreme events. To ensure the optimized initial conditions lead to physically meaningful and surface-relevant extreme temperatures across the entire domain, we analyze in App. C the 850-hPa temperature ($T_{850}$).
@@
To evaluate the robustness and physical realism of the NeuralGCM-optimized initial conditions, these perturbations should be tested in a conventional numerical weather prediction model (e.g., Environment and Climate Change Canada’s model buehner_implementation_2015). Such cross-model validation would reveal the universality of the results and help isolate NeuralGCM-specific biases. Additionally, running these scenarios in a fully physical model would explicitly account for land–atmosphere interactions and feedbacks, and assess whether the extreme trajectories persist under more detailed dynamics and physics.
Our method focuses on optimizing initial conditions, assuming the underlying model physics (whether learned or explicit) are fixed and skillful. An alternative approach could involve optimizing model parameters themselves (as done for example by alet2025skillfuljointprobabilisticweather to generate ensembles), though this would require careful regularization to ensure the resulting model remains physically plausible.
@@
The computational efficiency of the ML and hybrid models coupled with their differentiable properties, opens avenues for exploring extreme events—from heatwaves to precipitation extremes and compound disasters. For example, a similar optimization problem could be formulated for StormCast (pathak2024kilometerscaleconvectionallowingmodel) to allow us to search for extreme precipitation events in an emulated regional climate model. One could also formulate the loss function such that the large deviation theory rate function is minimized, leading to “typical” trajectories of extremes (10.1063/1.5084025; PhysRevX.13.041044). We could also envision loss functions with hard constraints on the perturbation which impose conservation laws as opposed to simply imposing small perturbations.
@@
We introduce a differentiable-storyline framework that leverages automatic differentiation in hybrid climate models to directly optimize initial conditions and generate physically coherent extreme-heatwave trajectories at a fraction of the computational cost of traditional ensemble methods. For example, our 50-step optimization run produced a more extreme event than any member of a 75-member ensemble, while using 33% less computational resources than it took to generate that ensemble. When applied to the PN2021 heatwave, our approach produces intensifications of nearly $3.7^{\circ}$C by isolating high-impact circulation patterns, specifically enhanced blocking and Rossby-wave, demonstrating both its dynamical fidelity and efficiency. While this proof-of-concept focuses on NeuralGCM and a single case study, the optimization paradigm is agnostic to model architecture and event type, offering a transformative tool for rapid, process-based risk assessment of diverse climate extremes in a warming world.
@@
Data availability: All data generated using the code and ERA5 data used in this study is openly available from the Copernicus Data Store (https://doi.org/10.24381/cds.143582cf) (ERA5).
Appendix A PN2021 ERA5
For reference, Figure 7 presents the atmospheric fields observed during the 2021 heatwave from the ERA5 reanalysis dataset. The figure displays the 1000-hPa temperature ($T_{1000}$) and the 500-hPa geopotential height ($Z_{500}$), characterizing the event. This illustrates the large-scale structure, such as the high-amplitude ridge in $Z_{500}$, associated with the event’s formation. Fig. 8 shows the other variables as in Fig. 5. The specific humidity, and winds are within the envelop of the NeuralGCM ensemble. The surface pressure on the other hand has a positive bias in NeuralGCM likely due to the representation of the surface in the coarse NeuralGCM model.
Figure 7: (a) Time series of 1000-hPa temperature for ERA5 for the PN2021. (b) Spatial map of average temperature anomalies from the ERA5 data. (c) Time series of 500-hPa geopotential height for ERA5. (d) Spatial map of 500-hPa geopotential height anomalies during the event period.
Figure 8: Equivalent of figure 5 using ERA5 data. Time series of (a) surface pressure (b) near-surface specific humidity (c) U-component of wind and (d) V-component of wind at 1000-hPa.
Appendix B Perturbation Spectrum
The perturbations introduced in the optimized runs are fully three-dimensional and span all horizontal and vertical levels. Due to this high dimensionality, it is challenging to visualize their full structure directly. To provide some insight into their characteristics, we show the spatial spectrum of the perturbations at selected vertical levels in Figure 9. This representation highlights the dominant spatial scales of the perturbations across the domain. A more detailed analysis of their full spatial structure could be informative, but is beyond the scope of this work.
Figure 9: Spectrum of the three-dimensional perturbations at selected vertical levels for the geopotential, temperature, kinetic energy and specific humidity.
Appendix C Temperature at 850-hPa
NeuralGCM utilizes $\sigma$-coordinates. Over regions with significant elevation, such as the Canadian Rockies, the 1000-hPa geopotential surface is often below ground level. Using the 1000-hPa temperature ($T_{1000}$) can therefore yield physically inconsistent values when optimizing for near-surface extreme events. To ensure the optimized initial conditions lead to physically meaningful and surface-relevant extreme temperatures across the entire domain, we analyze the 850-hPa temperature ($T_{850}$).
Fig. 10 presents the $T_{850}$ fields for the optimized extreme events. Despite the optimization targeting $T_{1000}$ the temperature at the 850-hPa level still exhibits a clear increase, exceeding the values observed in the ensemble simulations. We note that the magnitude of the anomaly found at 850-hPa is smaller than the maximum value achieved at the 1000-hPa level.
Figure 10: (a) Time series of 850-hPa temperature for EXP50, EXP75 and the ensemble. (b) Spatial map of average temperature anomalies from the EXP75 data.